# Learning High-Order Relationships of Brain Regions

### 02 Dec 2023 v1

@@ Abstract
Discovering reliable and informative interactions among brain regions from functional magnetic resonance imaging (fMRI) signals is essential in neuroscientific predictions of cognition. Most of the current methods fail to accurately characterize those interactions because they only focus on pairwise connections and overlook the high-order relationships of brain regions. We delve into this problem and argue that these high-order relationships should be maximally informative and minimally redundant (MIMR). However, identifying such high-order relationships is challenging and highly under-explored. Methods that can be tailored to our context are also non-existent. In response to this gap, we propose a novel method named HyBRiD that aims to extract MIMR high-order relationships from fMRI data. HyBRiD employs a Constructor to identify hyperedge structures, and a Weighter to compute a weight for each hyperedge. HyBRiD achieves the MIMR objective through an innovative information bottleneck framework named multi-head drop-bottleneck with theoretical guarantees. Our comprehensive experiments demonstrate the effectiveness of our model. Our model outperforms the state-of-the-art predictive model by an average of $12.1\%$, regarding the quality of hyperedges measured by CPM, a standard protocol for studying brain connections.
@@ 1 Introduction
Discovering reliable and informative relations among brain regions using fMRI signals is crucial for predicting human cognition in and understanding brain functions (Kucian et al., 2008; 2006; Li et al., 2015b; a; Satterthwaite et al., 2015; Wang et al., 2016).
@@
However, despite the clear multiplexity of the brain’s involvement in cognition (Logue & Gould, 2014; Barrasso-Catanzaro & Eslinger, 2016; Knauff & Wolf, 2010; Reineberg et al., 2022), the current imaging biomarker detection methods (Shen et al., 2017; Gao et al., 2019; Li et al., 2021) focus only on the contributing roles of the pairwise connectivity edges. In contrast, most brain functions involve distributed patterns of interactions among multiple regions (Semedo et al., 2019). For instance, executive planning requires the appropriate communication of signals across many distinct cortical areas (Logue & Gould, 2014). Pairwise connectivity, without considering the brain’s high-order structure, can lead to inconsistent findings across studies with low prediction performance. Although there are few works (Zu et al., 2016; Xiao et al., 2019; Li et al., 2022) working on discovering the high-order relationships of brain regions, they cannot effectively extract meaningful patterns towards a neurological outcome since 1) the discovering process of high-order relationships is not guided by cognitive targets; 2) the weak expressiveness of traditional machine learning methods (e.g. lasso regression) hinders the accurate modeling of brain region interactions (Cao et al., 2022; Richards et al., 2019).
@@
In this paper, we aim to identify high-order relationships that are informative toward a neurological outcome. However, unlike pairwise relations, the number of possible high-order relations is exponential. To identify the most informative ones from the exponential space, we propose our objective: maximally informative and minimally redundant (MIMR). That is to say, we aim to maximize the information contained in the high-order relationships towards a neurological outcome (informativeness) while diminishing the participation of unrelated brain regions (redundant). Such a criterion, on the one hand, ensures the predictive performance of these high-order relationships; on the other hand, it endows the model with the capacity to identify more succinct and interpretable structures (Yu et al., 2020; Miao et al., 2022a; b; Chen & Ying, 2023). A formal definition of the MIMR criterion could be found in Equation 8 from an information bottleneck view.
@@
Figure 1: We identify high-order relationships of brain regions, where hyperedges possess strong relevance to cognition (maximal informativeness). Meanwhile, they contain the least irrelevant information.
@@
We formulate high-order relationships as hyperedges in a hypergraph, where regions are treated as nodes. Unlike a traditional graph where edges connect only two nodes, a hypergraph allows edges, known as hyperedges, to connect any number of nodes. The hypergraph should be weighted, and the weights of hyperedges are considered as strengths of high-order relationships, which contain the information relevant to cognition (Figure 1).
@@
However, current methods for hypergraph construction, which mostly are based on neighbor distances and neighbor reconstruction (Wang et al., 2015; Liu et al., 2017; Jin et al., 2019; Huang et al., 2009), are unsuitable in our context for several reasons: 1) they are unable to learn MIMR hyperedges due to the absence of a tractable objective for learning such hyperedges. 2) they fall short of learning consistent structures across subjects, which contradicts the belief that the cognitive function mechanism of healthy human beings should be similar (Wang et al., 2023). 3) the number of hyperedges is restricted to the number of nodes, which may lead to sub-optimal performance. Furthermore, although information bottleneck (IB) is a prevailing solution to learn MIMR representations in deep learning (Kim et al., 2021; Alemi et al., 2016; Luo et al., 2019), existing IB methods focus on extracting compressed representations of inputs instead of identifying underlying structures such as hypergraphs. Harnessing the IB framework for identifying hypergraphs necessitates both architectural innovations and theoretical derivations.
@@
HyBRiD gets rid of searching in an exponential space through learning masks to identify hyperedges, which guarantees computational efficiency. Its feature-agnostic masking mechanism design ensures HyBRiD to learn consistent structures across subjects. Moreover, the model is equipped with a number of parallel heads, and each head is dedicated to a hyperedge. Through this, HyBRiD is able to identify any number of hyperedges, depending on how many heads it is equipped with. Additionally, the proposed multi-head drop-bottleneck theoretically guarantees the maximal informativeness and minimal redundancy of the identified hyperedges.
We evaluate our methods on $8$ datasets of human subjects under different conditions and at different participation times. We quantitatively evaluate our approach by a commonly used protocol for studying brain connections, CPM (Shen et al., 2017) (Appendix B), and show that our model outperforms the state-of-the-art deep learning models by an average of $12.1\%$ on a comprehensive benchmark. Our post-hoc analysis demonstrates that hyperedges of higher degrees are considered more significant, which indicates the significance of high-order relationships in human brains.
@@ 2 Problem Definition & Notations
We consider ABCD, which is a widely used benchmark in fMRI studies (5.1). This dataset is a collection of instances, where each instance corresponds to an individual human subject. An instance is represented by the pair $(X,Y)$. $X\in\mathbb{R}^{N\times d}$ represents the features for each subject, where $N$ is the number of brain regions and $d$ is the number of feature dimensions. The features are obtained by connectivity, consistent with previous works (5.1). $Y\in\mathbb{R}$ denotes the prediction target, such as fluid intelligence. Section 5 elaborates more details about datasets and preprocessing procedures. For how to obtain features of regions from raw data, see Paragraph 5.1 for more details.
@@
Based on the input $X$, HyBRiD aims to learn a weighted hypergraph of the brain, where regions are nodes. To achieve this, HyBRiD identifies a collection of hyperedges $H=(\bm{h}^{1},\bm{h}^{2},\cdots,\bm{h}^{K})$, and assigns weights $\bm{w}=[w^{1},w^{2},\cdots,w^{K}]^{T}$ for all hyperedges. These hyperedges and their weights, which represent strengths of hyperedges, are expected to be the most informative (i.e. relevant to the target $Y$) yet the least redundant.
@@
As mentioned before, we use $H$ to denote the collection of hyperedge structures and $\bm{h}^{k}$ to denote the $k$-th hyperedge. To associate hyperedges with node memberships, we use the following representation for a hyperedge:
$\bm{h}^{k}=\bm{m}^{k}\odot X\in\mathbb{R}^{N\times d}$ (1)
where $\bm{m}^{k}\in\{0,1\}^{N}$ is a mask vector and $\odot$ denotes broadcasting element-wise multiplication. In other words, each $\bm{h}^{k}$ is a random row-zeroed version of $X$.
@@ 3 Related Work
Existing hypergraph construction methods are mostly based on neighbor reconstruction and neighbor distances. For example, the $k$ nearest neighbor-based method (Huang et al., 2009) connects a centroid node and its $k$ nearest neighbors in the feature space to form a hyperedge. Wang et al. (2015); Liu et al. (2017); Jin et al. (2019); Xiao et al. (2019) further refine these neighbor connections through various regularization. However, the number of hyperedges of these methods is restricted to the number of nodes, and hyperedges obtained by these methods are inconsistent across instances. Zhang et al. (2022; 2018) proposed to iteratively refine a noisy hypergraph, which is obtained by the aforementioned methods. Therefore, they share the same limitations as the aforementioned methods. In addition, these methods are unable to learn MIMR hyperedges due to the absence of a tractable objective. Other methods such as attributed-based methods (Huang et al., 2015; Joslyn et al., 2019) are ill-adapted to our context since they require discrete labels or a prior graph topology. Different from these methods, we provide a way to learn a consistent hypergraph through a deep-learning model without any prior topology. Furthermore, thanks to the proposed multi-head drop-bottleneck, these hyperedges are theoretically ensured MIMR.
@@
Although some methods are working on high-order relationships in fMRI, they are limited or inconsistent with our MIMR objective. Xiao et al. (2019); Li et al. (2022) used the existing non-learning-based hypergraph construction methods, which may lead to noisy and inexpressive hypergraphs. Zu et al. (2016); Santoro et al. (2023) enumerated all hyperedges with degrees lower than 3, which can only discover a tiny portion of all possible hyperedges in exponential space and is not scalable to a large degree. Rosas et al. (2019) proposed O-information, which reflects the balance between redundancy and synergy. The O-information metric is utilized by Varley et al. (2023) to study fMRI data. However, the objective of these methods is not consistent with ours: although both of us are quantifying the redundancy of high-order relations, our method is to learn those that are most informative toward a cognition score, while theirs is to depict the synergy and redundancy within a system.
@@
Information bottleneck (IB) (Tishby et al., 2000) is originally a technique in data compression. The key idea is to extract a summary of data, which contains the most relevant information to the objective. Alemi et al. (2016) first employed an IB view of deep learning. After that, IB has been widely used in deep learning. The applications span areas such as computer vision (Luo et al., 2019; Peng et al., 2018), reinforcement learning (Goyal et al., 2019; Igl et al., 2019), natural language processing (Wang et al., 2020) and graph learning (Yu et al., 2020; 2022; Xu et al., 2021; Wu et al., 2020). Unlike these studies that use IB to extract a compressed representation or a select set of features, our approach focuses on identifying the underlying structures of the data.
Connectivity-based cognition prediction
Recently, deep learning techniques have been increasingly employed in predicting cognition based on the connectivity of brain regions. A substantial portion of these studies (Ahmedt-Aristizabal et al., 2021; Li et al., 2019; Cui et al., 2022b; Kan et al., 2022a; Cui et al., 2022a; Said et al., 2023) model the brain network as a graph, in which regions act as nodes and pairwise correlations form the edges. These methods predominantly utilize Graph Neural Networks (GNNs) to capture the connectivity information for predictions. In addition to GNNs, Kan et al. (2022b) proposed to use transformers with a specially designed readout module, leveraging multi-head attention mechanisms to capture pairwise connectivity. However, these methods heavily rely on pairwise connectivity and neglect more intricate higher-order relationships. This oversight, on the one hand, leads to sub-optimal prediction performances; and on the other hand, prevents domain experts from acquiring insightful neuroscience interpretations, given that human cognition often involves multiple regions.
@@ 4 Method
HyBRiD consists of a Constructor $\mathcal{F}_{c}$, a Weighter $\mathcal{F}_{w}$, and a LinearHead $\mathcal{F}_{l}$. At a high level, the Constructor $\mathcal{F}_{c}$ is responsible for identifying hyperedges $H$ from the data to construct the hypergraph. After that, the Weighter $\mathcal{F}_{w}$ calculates a weight for each hyperedge. Finally, based on all the weights $\bm{w}$, the LinearHead $\mathcal{F}_{l}$ predicts the label $Y$. An illustration of this pipeline is shown in Figure 2. The pipeline can be formulated as
@@
We will elaborate on all of the details of the architecture below.
@@ 4.1 Learning the hypergraph by Multi-head Masking
Given an instance, represented by $X=[X_{1},X_{2},\cdots,X_{N}]^{T}\in\mathbb{R}^{N\times d}$, where $X_{i}\in\mathbb{R^{d}}$ is a column vector representing the features of region $i$. These regions are nodes in the hypergraph we are going to construct. Hyperedges in the hypergraph can be beneficial in the learning task because it is essential to model the interactions between more than two regions.
@@
In this paragraph, we elaborate how the Constructor identifies the hyperedges, i.e. $H=\mathcal{F}_{c}(X)$.
@@
Specifically, to construct the $k$-th hyperedge, the Constructor’s $k$-th head outputs a column vector $\bm{m}^{k}\in\{0,1\}^{N}$, where each element in the vector corresponds to a region.
$\bm{m}^{k}$ $=[\mathbbm{1}(p^{k}_{\theta,1}),\mathbbm{1}(p^{k}_{\theta,2}),\cdots,\mathbbm{1}(p^{k}_{\theta,N})]^{T}\in\{0,1\}^{N}$ (3)
where $p^{k}_{\theta,i},i=1,2,\cdots,N$ are learnable parameters. $\mathbbm{1}:[0,1]\mapsto\{0,1\}$ is an indicator function, which is defined as $\mathbbm{1}(x)=1$ if $x>0.5$ and $\mathbbm{1}(x)=0$ if $x\leq 0.5$. And $\bm{m}^{k}$ is a column vector corresponding to the $k$-th hyperedge. Note that since there is no gradient defined for the indicator operation, we employ the stopgradient technique (Oord et al., 2017; Bengio et al., 2013) to approximate the gradient.
Figure 3: Architecture details of the Constructor and the Weighter when the number of nodes $N=6$ and the number of hyperedges $K=3$. At a high level, the Constructor learns the hyperedge structure by masking nodes. The Weighter computes the weight of each hyperedge based on the the hyperedge’s member nodes and their features.
In the vector $\bm{m}^{k}$, 00 indicates nodes that are masked out, and $1$ indicates nodes that are not masked. Nodes that are not masked are considered to form a hyperedge together. We use $\bm{h}^{k}$ to represent the masked version of $X$
$\bm{h}^{k}$ $=\bm{m}^{k}\odot X$ (4)
$=[\bm{m}^{k}_{1}X_{1},\bm{m}^{k}_{2}X_{2},\cdots,\bm{m}^{k}_{N}X_{N}]\in\mathbb{R}^{N\times d}$
where $\odot$ is the broadcast element-wise multiplication. $\bm{m}^{k}_{j}$ is the $j$-th element of the vector $\bm{m}^{k}$.
@@
After obtaining the structure (i.e. member nodes) of each hyperedge, the Weighter will calculate each hyperedge’s weight, which is supposed to indicate the importance of that hyperedge, based on the member nodes and their features, i.e. $\bm{w}=\mathcal{F}_{w}(H)$.
@@
$w^{k}=\operatorname{Readout}(\bm{h}^{k})=\operatorname{DimReduction}({\bm{m}^{k}}^{T}\bm{h}^{k})\in\mathbb{R}$ (6)
where $w^{k}$ is the weight of the $k$-th hyperedge and $\operatorname{DimReduction}$ is an MLP with ReLU activations, where the output dimension is 1. For all hyperedges, we obtain $K$ hyperedges in total, $\bm{w}=[w^{1},w^{2},\cdots,w^{K}]^{T}\in\mathbb{R}^{K}$. Finally, these weights will be fed into the final linear head to predict the label of the instance,
$\hat{Y}=\mathcal{F}_{l}(\bm{w})\in\mathbb{R}$ (7)
In contrast to previous hypergraph construction methods (Jin et al., 2019; Xiao et al., 2019), which identify hyperedges by refining neighbors and assign weights by aggregating node features, HyBRiD makes these procedures learnable and thus is able to identify MIMR hyperedges in a data-driven way through expressive neural networks. We decide the number of hyperedges $K$ according to the study in Appendix E.2.

### 13 Feb 2024 v2

@@ Abstract
Discovering reliable and informative relationships among brain regions from functional magnetic resonance imaging (fMRI) signals is essential in phenotypic predictions. Most of the current methods fail to accurately characterize those interactions because they only focus on pairwise connections and overlook the high-order relationships of brain regions. We propose that these high-order relationships should be maximally informative and minimally redundant (MIMR). However, identifying such high-order relationships is challenging and under-explored due to the exponential search space and the absence of a tractable objective. In response to this gap, we propose a novel method named HyBRiD which aims to extract MIMR high-order relationships from fMRI data. HyBRiD employs a Constructor to identify hyperedge structures, and a Weighter to compute a weight for each hyperedge, which avoids searching in exponential space. HyBRiD achieves the MIMR objective through an innovative information bottleneck framework named multi-head drop-bottleneck with theoretical guarantees. Our comprehensive experiments demonstrate the effectiveness of our model. Our model outperforms the state-of-the-art predictive model by an average of $11.2\%$, regarding the quality of hyperedges measured by CPM, a standard protocol for studying brain connections.
Machine Learning, ICML
@@ 1 Introduction
Discovering relations among brain regions toward a specific phenotypic outcome from fMRI signals has been a crucial area in neuroimaging research. Reliable and informative relations help neuroscientists and clinical professionals to better understand brain functions, and thus improve clinical diagnosis and treatments (Kucian et al., 2008, 2006; Li et al., 2015b, a; Satterthwaite et al., 2015; Wang et al., 2016).
@@
However, despite the clear multiplexity of the brain’s involvement in cognition (Logue & Gould, 2014; Barrasso-Catanzaro & Eslinger, 2016; Knauff & Wolf, 2010; Reineberg et al., 2022), current imaging biomarker detection methods (Shen et al., 2017; Gao et al., 2019; Li et al., 2021) focus only on the contributing roles of the pairwise connectivity edges. In contrast, most brain functions involve distributed patterns of interactions among multiple regions (Semedo et al., 2019). For instance, executive planning requires the appropriate communication of signals across many distinct cortical areas (Logue & Gould, 2014). Pairwise connectivity, without considering the brain’s high-order structure, can lead to inconsistent findings across studies with low prediction performance. Although recently there have been a few works (Zu et al., 2016; Xiao et al., 2019; Li et al., 2022) working on discovering the high-order relationships of brain regions, they are unable to effectively extract meaningful patterns. This is because these methods usually first identify some candidates of high-order relationships and then perform feature selection. The candidates are obtained by enumerating only a small portion of all possible relations, or by clustering methods that are unrelated to the target. As a result, most informative high-order relations are likely excluded at the first stage. This inspires us to solve the problem in an end-to-end manner through a more expressive model.
@@
In this paper, we aim to identify high-order relationships that are informative towards a phenotypic outcome, such as a cognition score. However, unlike pairwise relations, the number of possible high-order relations is exponential. To identify the most informative ones from the exponential space, we propose our objective: maximally informative and minimally redundant (MIMR). That is to say, we maximize the information contained in the high-order relationships towards a neurological outcome (informativeness) while diminishing the participation of unrelated brain regions (redundant). Such a criterion, on the one hand, ensures the predictive performance of these high-order relationships; on the other hand, it endows the model with the capacity to identify more succinct and interpretable structures (Yu et al., 2020; Miao et al., 2022a, b; Chen & Ying, 2023). A formal definition of the MIMR criterion could be found in Equation 8 from an information bottleneck point of view.
@@
Figure 1: We identify high-order relationships of brain regions, where hyperedge structures and weights possess strong relevance to cognition (maximal informativeness). Meanwhile, they contain the least irrelevant information.
@@
We formulate high-order relationships as weighted hyperedges in a hypergraph, where regions are treated as nodes. Unlike a traditional graph where edges connect only two nodes, a hypergraph allows edges, known as hyperedges, to connect any number of nodes. The hypergraph should be weighted, and the weights of hyperedges are considered as strengths of high-order relationships, which contain the information relevant to the outcome (Figure 1).
@@
However, current methods for hypergraph construction, which are mostly based on neighbor distances and neighbor reconstruction (Wang et al., 2015; Liu et al., 2017; Jin et al., 2019; Huang et al., 2009), are unsuitable in our context for several reasons: 1) they are unable to learn MIMR hyperedges due to the absence of a tractable objective for learning such hyperedges. 2) they fall short of learning consistent structures across subjects, which contradicts the belief that the cognitive function mechanism of healthy human beings should be similar (Wang et al., 2023). 3) the number of hyperedges is restricted to the number of nodes, which may lead to sub-optimal performance. Furthermore, although information bottleneck (IB) has been a prevailing solution to learn MIMR representations in deep learning (Kim et al., 2021; Alemi et al., 2016; Luo et al., 2019), existing IB methods focus on extracting compressed representations of inputs instead of identifying underlying structures such as hypergraphs. Harnessing the IB framework for identifying hypergraphs necessitates both architectural innovations and theoretical derivations.
@@
HyBRiD avoids searching in an exponential space through learning masks to identify hyperedges, which guarantees efficiency. Its feature-agnostic masking mechanism design ensures HyBRiD to learn consistent structures across subjects. Moreover, the model is equipped with a number of parallel heads, each of which is dedicated to a hyperedge. Through this, HyBRiD is able to identify any number of hyperedges, depending on how many heads it is equipped with. Additionally, the proposed multi-head drop-bottleneck theoretically guarantees the maximal informativeness and minimal redundancy of the identified hyperedges.
We evaluate our methods on the open-source ABIDE dataset and the restricted ABCD dataset. We quantitatively evaluate our approach by a commonly used protocol for studying brain connections, CPM (Shen et al., 2017) (Appendix B), and show that our model outperforms the state-of-the-art deep learning models by an average of $11.2\%$ on a comprehensive benchmark. Our post-hoc analysis demonstrates that hyperedges of higher degrees are considered more significant, which indicates the significance of high-order relationships in human brains.
@@ 2 Problem Definition & Notations
Our dataset is a collection of human subject’s features and their phenotypic outcomes, which is represented by the pair $(X,Y)$ for each subject. $X\in\mathbb{R}^{N\times d}$ represents the features for each subject, where $N$ is the number of brain regions and $d$ is the feature size. Consistent with previous works (Kan et al., 2022b; Li et al., 2021), the features are Pearson correlations derived from fMRI time series. $Y\in\mathbb{R}$ denotes the phenotypic outcome, such as the intelligent quotient. Section 5 and Appendix D will elaborate more details about datasets and preprocessing procedures.
@@
Based on the input $X$, HyBRiD aims to learn a weighted hypergraph of the brain, where regions are nodes. To achieve this, HyBRiD identifies a collection of hyperedges $H=(\boldsymbol{h}^{1},\boldsymbol{h}^{2},\cdots,\boldsymbol{h}^{K})$, and assigns weights $\boldsymbol{w}=[w^{1},w^{2},\cdots,w^{K}]^{T}$ for all hyperedges. These hyperedges and their weights, which represent strengths of hyperedges, are expected to be the most informative towards $Y$ yet the least redundant.
@@
As mentioned before, we use $H$ to denote the collection of hyperedge structures and $\boldsymbol{h}^{k}$ to denote the $k$-th hyperedge. To associate hyperedges with node memberships, we use the following representation for a hyperedge:
$\boldsymbol{h}^{k}=\boldsymbol{m}^{k}\odot X\in\mathbb{R}^{N\times d},$ (1)
where $\boldsymbol{m}^{k}\in\{0,1\}^{N}$ is a mask vector and $\odot$ denotes broadcasting element-wise multiplication. In other words, each $\boldsymbol{h}^{k}$ is a randomly row-zeroed version of $X$.
@@ 3 Related Work
Existing hypergraph construction methods are mostly based on neighbor reconstruction and neighbor distances. For example, the $k$ nearest neighbor-based method (Huang et al., 2009) connects a centroid node and its $k$ nearest neighbors in the feature space to form a hyperedge. Wang et al. (2015); Liu et al. (2017); Jin et al. (2019); Xiao et al. (2019) further refine these neighbor connections through various regularization. However, the number of hyperedges of these methods is restricted to the number of nodes, and hyperedges obtained by these methods are inconsistent across instances. Zhang et al. (2022, 2018) proposed to iteratively refine a noisy hypergraph, which is obtained by the aforementioned methods. Therefore, they share the same limitations as the aforementioned methods. In addition, these methods are unable to learn MIMR hyperedges due to the absence of a tractable objective. Other methods, such as attributed-based methods (Huang et al., 2015; Joslyn et al., 2019), are ill-adapted to our context since they require discrete labels. Different from these methods, we provide a way to learn a consistent hypergraph through a deep-learning model. Furthermore, thanks to the proposed multi-head drop-bottleneck, these hyperedges are theoretically ensured MIMR.
@@
Although there are some methods working on high-order relationships in fMRI, they have limitations and are inconsistent with our MIMR objective. Xiao et al. (2019); Li et al. (2022) used the existing non-learning-based hypergraph construction methods, which may lead to noisy and inexpressive hypergraphs. Zu et al. (2016); Santoro et al. (2023) enumerated all hyperedges with degrees lower than 3, which can only discover a tiny portion of all possible hyperedges in exponential space and is not scalable to a large degree. Rosas et al. (2019) proposed O-information, which reflects the balance between redundancy and synergy. The O-information metric is utilized by Varley et al. (2023) to study fMRI data. However, the objective of these methods is not consistent with ours: although both of us are quantifying the redundancy of high-order relations, our method is to learn those that are most informative toward a cognition score, while theirs is to depict the synergy and redundancy within a system.
@@
Information bottleneck (IB) (Tishby et al., 2000) is originally a technique in data compression. The key idea is to extract a summary of data, which contains the most relevant information to the objective. Alemi et al. (2016) first employed an IB view of deep learning. After that, IB has been widely used in deep learning. The applications span areas such as computer vision (Luo et al., 2019; Peng et al., 2018), reinforcement learning (Goyal et al., 2019; Igl et al., 2019), natural language processing (Wang et al., 2020) and graph learning (Yu et al., 2020, 2022; Xu et al., 2021; Wu et al., 2020). Unlike these studies that use IB to extract a compressed representation or a select set of features, our approach focuses on identifying the underlying structures of the data.
Connectivity-based Phenotypic Prediction
Recently, deep learning techniques have been increasingly employed in predicting phenotypic outcomes based on the connectivity of brain regions. Most works (Ahmedt-Aristizabal et al., 2021; Li et al., 2019; Cui et al., 2022b; Kan et al., 2022a; Cui et al., 2022a; Said et al., 2023) model the brain network as a graph, in which regions act as nodes and pairwise correlations form the edges. These methods predominantly utilize Graph Neural Networks (GNNs) to capture the connectivity information for predictions. In addition to GNNs, Kan et al. (2022b) proposed to use transformers with a specially designed readout module, leveraging multi-head attention mechanisms to capture pairwise connectivity. However, all of these methods heavily rely on pairwise connectivity and neglect more intricate higher-order relationships. This oversight, on the one hand, leads to sub-optimal prediction performances and, on the other hand, prevents domain experts from acquiring insightful neuroscience interpretations, given that brain functions often involves multiple regions.
@@ 4 Method
HyBRiD consists of a Constructor $\mathcal{F}_{c}$, a Weighter $\mathcal{F}_{w}$, and a LinearHead $\mathcal{F}_{l}$. At a high level, the Constructor $\mathcal{F}_{c}$ is responsible for identifying hyperedges $H$ from the data to construct the hypergraph. After that, the Weighter $\mathcal{F}_{w}$ calculates a weight for each hyperedge. Finally, based on all the weights $\boldsymbol{w}$, the LinearHead $\mathcal{F}_{l}$ predicts the label $Y$. An illustration of this pipeline is shown in Figure 2. The pipeline can be formulated as
@@
We will elaborate on the details of the architecture below.
@@ 4.1 Learning the Hypergraph by Multi-head Masking
Each instance (human subject) is represented by $X=[X_{1},X_{2},\cdots,X_{N}]^{T}\in\mathbb{R}^{N\times d}$, where $X_{i}\in\mathbb{R^{d}}$ is a column vector representing the features of region $i$. These regions are nodes in the hypergraph we are going to construct. Hyperedges in the hypergraph can be beneficial in the learning process below because it is essential to model the relationships between more than two regions.
@@
In this paragraph, we elaborate on how the Constructor identifies the hyperedges, i.e. $H=\mathcal{F}_{c}(X)$.
@@
Specifically, to construct the $k$-th hyperedge, the Constructor’s $k$-th head outputs a column vector $\boldsymbol{m}^{k}\in\{0,1\}^{N}$, where each element in the vector corresponds to a brain region,
$\boldsymbol{m}^{k}=[\mathbbm{1}(p^{k}_{\theta,1}),\mathbbm{1}(p^{k}_{\theta,2}),\cdots,\mathbbm{1}(p^{k}_{\theta,N})]^{T}\in\{0,1\}^{N},$ (3)
where $p^{k}_{\theta,i}\in[0,1],i=1,2,\cdots,N$ are learnable probabilities. $\mathbbm{1}:[0,1]\mapsto\{0,1\}$ is an indicator function, which is defined as $\mathbbm{1}(x)=1$ if $x>0.5$ and $\mathbbm{1}(x)=0$ if $x\leq 0.5$. And $\boldsymbol{m}^{k}$ is a column vector corresponding to the $k$-th hyperedge. Note that since there is no gradient defined for the indicator operation, we employ the stop-gradient technique (Oord et al., 2017; Bengio et al., 2013) to approximate it.
Figure 3: Architecture details of the Constructor and the Weighter when the number of nodes $N=6$ and the number of hyperedges $K=3$. At a high level, the Constructor learns the hyperedge structure by masking nodes. The Weighter computes the weight of each hyperedge based on the hyperedge’s member nodes and their features.
In the vector $\boldsymbol{m}^{k}$, 00 indicates nodes that are masked out, and $1$ indicates nodes that are not masked. Nodes that are not masked are considered to form a hyperedge together. We use $\boldsymbol{h}^{k}$ to represent the masked version of $X$
$\boldsymbol{h}^{k}$ $=\boldsymbol{m}^{k}\odot X$ (4)
$=[\boldsymbol{m}^{k}_{1}X_{1},\boldsymbol{m}^{k}_{2}X_{2},\cdots,\boldsymbol{m}^{k}_{N}X_{N}]\in\mathbb{R}^{N\times d},$
where $\odot$ is the broadcast element-wise multiplication. $\boldsymbol{m}^{k}_{j}$ is the $j$-th element of the vector $\boldsymbol{m}^{k}$.
@@
After obtaining the structure (i.e. member nodes) of each hyperedge, the Weighter will calculate each hyperedge’s weight, which is supposed to indicate the importance of that hyperedge, based on the member nodes and their features, i.e. $\boldsymbol{w}=\mathcal{F}_{w}(H)$.
@@
$w^{k}=\operatorname{Readout}(\boldsymbol{h}^{k})=\operatorname{DimReduction}({\boldsymbol{m}^{k}}^{T}\boldsymbol{h}^{k})\in\mathbb{R},$ (6)
where $w^{k}$ is the weight of the $k$-th hyperedge and $\operatorname{DimReduction}$ is an MLP with ReLU activations, where the output dimension is $1$. For all hyperedges, we obtain $K$ hyperedges in total, $\boldsymbol{w}=[w^{1},w^{2},\cdots,w^{K}]^{T}\in\mathbb{R}^{K}$. Finally, these weights will be fed into the final linear head to predict the label of the instance,
$\hat{Y}=\mathcal{F}_{l}(\boldsymbol{w})\in\mathbb{R}.$ (7)
In contrast to previous hypergraph construction methods (Jin et al., 2019; Xiao et al., 2019), which identify hyperedges by refining neighbors and simply aggregating node features, HyBRiD makes these procedures learnable and thus is able to identify MIMR hyperedges in a data-driven way through expressive neural networks. The number of hyperedges $K$ is decided according to the study in Appendix H.1.
Computational Complexity
The computational complexity of our model is $O(N^{2}K)$, which is just the same scale as that of MLPs even though we are addressing a more challenging task: identifying high-order relationships in an exponential space. Details of the complexity calculation can be found in Appendix C.